# R1-Zero’s “Aha Moment” in Visual Reasoning on a 2B Non-SFT Model

### v1

@@ Abstract
Recently DeepSeek R1 demonstrated how reinforcement learning with simple rule-based incentives can enable autonomous development of complex reasoning in large language models, characterized by the "aha moment", in which the model manifest self-reflection and increased response length during training. However, attempts to extend this success to multimodal reasoning often failed to reproduce these key characteristics. In this report, we present the first successful replication of these emergent characteristics for multimodal reasoning on only a non-SFT 2B model. Starting with Qwen2-VL-2B and applying reinforcement learning directly on the SAT dataset, our model achieves 59.47% accuracy on CVBench, outperforming the base model by approximately ~30% and exceeding both SFT setting by ~2%. In addition, we share our failed attempts and insights in attempting to achieve R1-like reasoning using RL with instruct models. aiming to shed light on the challenges involved. Our key observations include: (1) applying RL on instruct model often results in trivial reasoning trajectories, and (2) naive length reward are ineffective in eliciting reasoning capabilities. The project code is available at https://github.com/turningpoint-ai/VisualThinker-R1-Zero
@@ 2.1 Multimodal Reasoning
Researchers have demonstrated that LLMs can be post-trained to elicit enhanced reasoning abilities [8]. With pre-trained visual encoders to understand visual content alongside textual data, multimodal LLM’s reasoning abilities are investigated and enhancement typically requires sophisticated prompting designs [6] or large amounts of reasoning training data [14; 18]. The research community is increasingly interested on developing more natural methods to incentivize higher intelligence in models without relying on extensive supervised data or complex prompting techniques.
@@ 3 VisualThinker R1 Zero
In this part, we demonstrate VisualThinker-R1-Zero have showed an emergence of “aha moment” by directly applying RL training to the non-SFT base model, leading to a superior performance on vision-centric tasks on non-SFT 2B models.
@@
Our method builds on Qwen-2-VL-2B [16] as the base model, applying GRPO [12] with a tailored chat template and prompting strategy to enhance its reasoning capabilities. We posit that applying GRPO to the base model may be a more efficient and effective way to replicate multimodal R1 reasoning compared to training instruct fine-tuned model. As we will show in Section 5, the instruction-tuned variant, Qwen-2-VL-2B-Instruct, demands significantly more training resources as a base model yet struggles to replicate the emergent "aha moment" observed in DeepSeek R1 with notable failure modes.

### v2

@@ Abstract
The recent DeepSeek-R1 demonstrated how reinforcement learning with simple rule-based reward can enable autonomous development of complex reasoning in large language models, characterized by the "aha moment", in which the model manifest self-reflection and increased response length during training. However, attempts to extend this success to multimodal reasoning often failed to reproduce these key characteristics. In this report, we present the first successful replication of these emergent characteristics for multimodal reasoning on only a non-SFT 2B model. Starting with Qwen2-VL-2B and applying reinforcement learning directly on the SAT dataset, our model achieves 59.47% accuracy on CVBench, outperforming the base model by approximately ~30% and exceeding both SFT setting by ~2%. In addition, we share our failed attempts and insights in attempting to achieve R1-like reasoning using RL with instruct models, aiming to shed light on the challenges involved. Our key observations include: (1) applying RL on instruct model often results in trivial reasoning trajectories, and (2) naive length reward are ineffective in eliciting reasoning capabilities. The project code is available at https://github.com/turningpoint-ai/VisualThinker-R1-Zero
@@ 2.1 Multimodal Reasoning
Researchers have demonstrated that LLMs can be post-trained to elicit enhanced reasoning abilities [8]. With pre-trained visual encoders to understand visual content alongside textual data, multimodal LLM’s reasoning abilities are investigated and enhancement typically requires sophisticated prompting designs [6] or large amounts of reasoning training data [14; 18]. The research community is increasingly interested in developing more natural methods to incentivize higher intelligence in models without relying on extensive supervised data or complex prompting techniques.
@@ 3 VisualThinker R1 Zero
In this section, we demonstrate VisualThinker-R1-Zero have showed an emergence of “aha moment” by directly applying RL training to the non-SFT base model, leading to a superior performance on vision-centric tasks on non-SFT 2B models.
@@
Our method builds on Qwen-2-VL-2B [16] as the base model, applying GRPO [12] with a tailored chat template and prompting strategy to enhance its reasoning capabilities. We posit that applying GRPO to the base model may be a more efficient and effective way to replicate multimodal R1 reasoning compared to training instruct fine-tuned model. As we will show in Section 5, the instruction-tuned variant, Qwen-2-VL-2B-Instruct, demands significantly more training resources as a base model yet struggles to replicate the emergent ”aha moment” observed in DeepSeek R1 with notable failure modes.